mathslover (mathslover):

Application of Derivatives Problem .... Find the intervals where the function \(f(x) = x^2 e^x\) is increasing and decreasing.

mathslover (mathslover):

@ganeshie8 and @BSwan This is what I have done yet. From the last problem as I differentiated it first, so, I will also differentiate this one. f'(x) = \(2xe^x + x^2 e^x\)

OpenStudy (anonymous):

good , then ?

mathslover (mathslover):

\(f'(x) = x e^x ( x + 2) \) ?

mathslover (mathslover):

For increasing , it is true that it should be \(\ge\) 0 and for decreasing it should be \(\le 0\) , right?

OpenStudy (anonymous):

ahaaa :D

mathslover (mathslover):

So, making conditions : For \(\bf{Increasing}\) : \(\boxed{\mathsf{xe^x (x+2) \ge 0 }}\) For \(\bf{Decreasing}\) : \(\boxed{\mathsf{xe^x(x+2) \le 0}}\)

mathslover (mathslover):

I'm not sure how to go on from here. Though, it should be easy to continue from this point, but am not sure how to do that.

ganeshie8 (ganeshie8):

e^x is always positive

mathslover (mathslover):

Oh... that may help.

mathslover (mathslover):

So, x can be lesser than -2 ... and x can be greater than 0 .. right?

mathslover (mathslover):

for increasing.

ganeshie8 (ganeshie8):

x(x+2) >=0 when x <-2 or x>0 x(x+2) <= 0 when x>-2 and x< 0

mathslover (mathslover):

\(x \in (-\infty , -2) \cup (0 , \infty)\)

mathslover (mathslover):

That was for increasing^ And for decreasing, I get : \( x \in (-2,0) \)

mathslover (mathslover):

o.O no...............!

OpenStudy (anonymous):

pick up any points in the intervals to see if it increas or decrease with f'(x) >0 test

mathslover (mathslover):

For increasing... \(x \in (-\infty , 2] \cup [ 0 ,\infty)\) and for decreasing.. \(x \in [-2,0] \) Is this right?

ganeshie8 (ganeshie8):

Looks good !

mathslover (mathslover):

I'm confused about the closed and open brackets. What the book (course) , says is : f(x) is increasing in \([-\infty , -2) \cup [0, \infty)\) f(x) is decreasing in \((-2,0]\)

ganeshie8 (ganeshie8):

it is illegal to include \([-\infty \)

mathslover (mathslover):

Yeah... Moreover, -2 should be included in increasing as well as in decreasing.

ganeshie8 (ganeshie8):

^Agree !

OpenStudy (anonymous):

i dnt dnt like the \(\cup\) notation xD

mathslover (mathslover):

Probably, just a printing mistake. As this is what it says. @BSwan -> I used to hate Cup and cap notations, but, now when I get them, I love them!

ganeshie8 (ganeshie8):

haha @BSwan please medal @mathslover :)

OpenStudy (anonymous):

well its not like that , previously i state why its not true to put them , so they still teach this in pre cal although it make a lote of doubts !

mathslover (mathslover):

:-) Okay, I have got some more question. I should close it now. Thanks a lot @BSwan and @ganeshie8 for being there to help me out everytime. I really thank you both for that.

OpenStudy (anonymous):

np :)